# Traffic Flow and Speed Monitoring Based On Optical Fiber Distributed Acoustic Sensor

### v1

@@ III-C2 Peaks location search algorithm
The vehicle position detection algorithm flow is shown in Algorithm 1.
@@ 3 $k_{s}$leftarrow FindPeak(ButterworthFilter($D_{1:m,1}$));
III-D Vehicle trajectory extraction
Figure 5: Workflow of the total algorithm.
In this section, we propose an algorithm for line-by-line matching based on the vehicle motion model. According to Newton’s first law, vehicles can typically be regarded as moving at a uniform velocity in the absence of lane changes, emergency stops, or overtaking maneuvers.
Considering the real highway scenario in this study, the space-time information reflected by optical fiber within the optical fiber waterfall diagram exhibits a straight-line pattern. This straight line’s slope is related to velocity, while the intercept is related to arrival time. This characteristic aligns well with a uniform linear motion model represented as:
$x=vt+x_{0}$ (10)
where $v$ is the vehicle’s velocity, $x_{0}$ is the initial position where the vehicle appears, and $x$ is the vehicle’s position at time $t$.
Figure 6: Trajectory extraction algorithm. The black line segment represents the velocity range and the red dot represents trajectory key points.
Therefore, after determining the starting position via vehicle position detection, knowing the vehicle’s velocity enables the determination of its position at different points along its trajectory. However, in the case of vehicle congestion, where each vehicle moves within a fixed range, there is a risk of multiple vehicles overlapping in their tracks. To address this challenge, we expand the speed range of each vehicle from a static line segment to a dynamic line segment. The selection of this dynamic line segment is determined by the slope of the line fitted using historical position and velocity, incorporating a certain degree of confidence. Fig. 6 illustrates the dynamic line segments within complex tunnel scenes, facilitating the tracking of key points in vehicle trajectories through line-by-line matching.
$y(x,\omega)=\omega_{0}+\omega_{1}x+\omega_{2}x+\ldots+w_{M}x^{M}=\sum_{j=0}^{M}\omega_{j}x^{j}$ (11)
$E(\omega)=\frac{1}{2}\sum_{n=1}^{N}(y(x_{n},\omega)-t_{n})^{2}$ (12)
The vehicle trajectory extraction algorithm flow is shown in Algorithm 2.
@@ III-D Vehicle trajectory extraction
1 Apply Algorithm 1 to $A=(a_{ij})$, returning preprocessed optical fiber matrix $D=(d_{ij})$; vehicle counts S and appearance time of each vehicle $k_{s}$. $(k_{s}+\alpha,l_{k_{s}+\alpha})$ is a pair of time and distance position points of each extracted vehicle trajectory;
@@
4 Select the initial velocity interval $v_{\min}^{(1)}$ to $v_{\max}^{(1)}$;
@@ 5 Calculate the initial distance interval $x_{\min}^{(1)}$ to $x_{\max}^{(1)}$;
6 foreach s = 1 to S; $\sigma^{(s)}\in\Sigma^{(S)}$ do
@@ 10             Add pair $(k_{s}+1,l_{k_{s}+1})$ to each vehicle trajectory point set $\sigma^{(s)}$;
13             break;
14
15foreach s = 1 to S; $\sigma^{(s)}\in\Sigma^{(S)}$ do
@@ 24                   Add pair $(k_{s}+\alpha,l_{k_{s}+\alpha})$ to each vehicle trajectory point set $\sigma^{(s)}$;
Based on the extraction results of key points for a single vehicle, we can deduce the vehicle’s approximate uniform-speed trajectory through polynomial fitting, as illustrated in Fig. 7.
The workflow diagram is shown in Fig. 5.
Figure 7: Trajectory extraction results. (a) trajectory key points of a large truck, and (b) trajectory fitting results of it.
IV Experiments
IV-A Datasets
In this study, we conducted data acquisition and analysis in two real highway scenarios, with a particular focus on a tunnel scenario. As our optical fiber is laid on only one side of the road, the target areas for this paper are explicitly stated. The algorithm proposed aims to extract vehicle trajectories within these target areas.
@@ IV-A Datasets
Target area in Jiurui tunnel dataset: The Jiujiang-Ruichang Expressway, situated in Jiujiang City, Jiangxi Province, China, represents a real highway tunnel scenario with a single carriageway featuring two lanes. Our target area focuses on the right lane, disregarding the left lane.
To assess the effectiveness of the optical fiber equipment, cameras were installed on the same right side of the road for data comparison. It is important to note that camera data is used here solely to evaluate the reliability of data collected by optical fiber. The algorithm proposed in this paper exclusively relies on optical fiber data and is independent of camera data.
IV-A1 Jingjintang Expressway Dataset
We selected a 320-meter section of Jingjintang Expressway for our study. Simultaneously, a camera mounted on the road captures movement information of the vehicles. Positioned at the starting point of the optical fiber laying side, the camera primarily monitors vehicles in the target area, as depicted in Fig. 8.
@@ IV-A1 Jingjintang Expressway Dataset
When a vehicle traverses the area, the Distributed Acoustic Sensing (DAS) system records the corresponding vibration signals. To present the vehicles’ trajectory more intuitively, we utilize waterfall diagrams to plot the collected signal values, as illustrated in Fig. 9.
Figure 9: Optical fiber waterfall diagrams of vehicles passing at intervals in Jingjintang expressway dataset. (a) three trucks pass at intervals, (b) three cars pass at intervals, and (c) one truck and four cars pass at intervals.
In the waterfall diagram, the light blue color represents background noise, while prominently colored lines indicate vehicles. The fiber stripe extending from the top left to the bottom right corresponds to vehicles within the target area, and the stripe extending from the top right to the bottom left represents vehicles on the opposite side. Notably, the vehicle’s trajectory is depicted as an approximately straight line within the diagram. This is attributed to the fact that most vehicles on the highway maintain high and consistent velocities when traveling in lanes with unchanged speeds. In this representation, the intercept of the straight line signifies the starting position of the vehicle, and the absolute value of the slope represents its velocity.
While optical fiber waterfall diagrams excel at representing real-time vehicle location data, applying them in actual highway scenarios reveals certain inevitable challenges, as depicted in Fig. 10.
@@
Figure 10: Displayed above are several representative waterfall diagrams featuring inaccurate optical fiber signals in the Jingjintang Expressway dataset. In the first row, optical fiber waterfall diagrams capture trucks passing nearly simultaneously in different lanes, significantly overlapping with neighboring cars. Moving to the second row, the diagrams illustrate vehicles from the opposite lane causing interference with the signals of vehicles in the target area. The presence of opposite vehicles is indicated by fiber stripes extending from the top right to the bottom left.
@@
(1) Owing to high sensitivity and rapid response characteristics of optical fiber, large trucks are prone to generating large signal amplitudes. In instances where a car passes concurrently in close proximity, the car’s signal may become entirely submerged beneath the truck’s signal.
@@
(2) Considering the relatively high number of large trucks on the highway, each truck generates substantial optical fiber signals. When multiple large vehicles traverse the roadway consecutively, they tend to interfere with each other, posing a challenge in distinguishing the count of vehicles within this cluster.
@@
(3) While the signals from vehicles traveling in the opposite lane are not as clear as those within the target area, signals from trucks in the opposing lane can still interfere with the traffic signal data we aim to capture.
@@
Additionally, we show a comparison between fiber stripes and camera ground truth to assess the accuracy of optical fiber. Fig. 11 presents the statistics of the number of trucks and cars for each minute within two hours. The full lines represent the ground truth detected by the camera, while the dashed lines represent the estimated counts detected by the optical fiber stripes. Comparing the results in Table. I, the optical fiber can capture nearly all the information about trucks. However, due to the aforementioned challenges leading to inaccurate optical fiber signals, the counting result for cars is lower than ground truth.
@@
TABLE I: Comparison of real traffic flow and optical fiber stripe counts within 2 hours of Jingjintang expressway dataset.
@@
Vehicle type Fiber stripes counts Camera counts Accuracy rate
@@
(a) Trucks.
(b) Cars.
Figure 11: Vehicle statistics counts in Jingjintang expressway dataset. The blue and orange lines show the counting results of the camera and the fiber stripe respectively.
IV-A2 Jiurui Tunnel Dataset
We also selected a test section in the Jiurui tunnel, including an 80-meter section and the entire 2000-meter tunnel.
Similarly, we draw optical fiber waterfall diagrams to get a more intuitive view of the vehicles in Fig. 12. Compared with Jingjintang expressway dataset, the optical fiber stripes in Jiurui tunnel dataset are more strong and obvious. This distinction arises from the tunnel’s configuration, consisting of just two lanes in the same direction, with trucks entering from the outer lane. Consequently, the distinct signals on the waterfall diagram correspond to trucks on the outer lane. Additionally, due to the attenuation effect of the tunnel, the signals from cars in the inner lane are too weak to be adequately recorded. Therefore, the relatively faint signals on the waterfall diagram represent cars in the inner lane.
@@ IV-A2 Jiurui Tunnel Dataset
The Jiurui tunnel dataset of short distance in Fig. 12 comprises 16 sets of original optical fiber data, with each set a collection time of 3-4 minutes and a collection distance of 80 meters. Among these, 7 datasets include corresponding video recordings. Comparing these optical fiber stripes in the optical fiber waterfall diagrams with the actual traffic flow in the videos in Fig. II, we observed that due to signal attenuation caused by the tunnel, the fiber stripes of small vehicles on the outer lane were faint, especially electric vehicles, which could not be adequately captured by the Distributed Acoustic Sensing (DAS) system. This explains the observed variance between optical fiber data acquisition and real traffic flow in the highway tunnel scenario.
TABLE II: Comparison of real traffic flow and optical fiber stripe counts in Jiurui tunnel dataset of short distance.
Collection time Fiber stripes counts Camera counts Accuracy rate
@@
In complex Jiurui tunnel long-distance scenarios shown in Fig. 2, prevalent conditions include low-speed limits and restrictions on lane changes and overtaking. Consequently, heavy traffic flows can lead to traffic congestion, as evident from the optical fiber waterfall diagrams. These diagrams reveal occurrences of multiple lines intersecting (congestion) and even overlapping (illegal overtaking) due to the specific traffic conditions and regulations enforced in the tunnel. The purpose of this paper is to track the whole trajectory of vehicles, even when the lane change problem is encountered.
IV-B Metrics

### v2

@@ 3.3 Vehicle Position Detection
We propose an algorithm for vehicle detection based on the Butterworth filter and peaks location search algorithm. The vehicle position detection algorithm flow is shown in Algorithm 1.
@@ 4 $k_{s}$leftarrow FindPeak(ButterworthFilter($D_{1:m,1}$));
3.3.1 Smoothing method of Butterworth Filter
For optical fiber waterfall diagrams after preprocessing, we further adopt Butterworth low-pass filtering on the signals at the starting position of optical fibers. Butterworth filter is a low-pass filter with maximum flat amplitude response in the pass-band [35]. Compared with analog filter, Butterworth filter has many advantages such as high precision, stability, flexibility, and no impedance matching [36]. Butterworth Filter has been widely used in many fields such as automatic control, image and communication. The Butterworth filter has two important parameters, order $N$ and normalized cut-off frequency $W_{n}$. The order determines the steepness of the frequency response curve, and the cut-off frequency is usually used to control the frequency characteristics. In optical fiber signal processing, noise mostly presents high frequency characteristics, so we choose Butterworth low-pass filter to remove the high frequency noise in the image and smooth signal details. In this way, we can further separate components of different frequencies, so as to make the signals even flatter.
3.3.2 Peaks location search algorithm
The features of wave peaks satisfy the principle that the first derivative is zero and the second derivative is negative, the position of the wave peaks can be searched as follows:
(1) Define the projection curve as $V=[v_{1},v_{2},\ldots,v_{n}]$.
(2) Calculate the first-order difference vector ${\rm{Diff}}_{v}(i)$ for $i\in 1,2,\ldots,N-1$:
${\rm{Diff}}_{v}(i)=v_{i+1}-v_{i}.$ (8)
(3) Apply the sign function to the difference vector ${S}(i)={\rm{sign}}({\rm{Diff}}_{v})$, where
(4) Iterate through $S(i)$ from the tail to perform the following:
${S}(i)=\begin{cases}1,\text{if }{S}(i)=0\text{ and }{S}(i+1)\geq 0;\\ -1,\text{if }{S}(i)=0\text{ and }{S}(i+1)<0.\end{cases}$ (10)
(5) Perform the first-order difference operation on $S(i)$ to obtain $R(i)={\rm{Diff}}_{v}(S(i))$.
(6) Traverse the difference vector $R$. If $R(i)=-2$, then $i+1$ is crest of the projection vector $V$, and the corresponding crest is $V(i+1)$. If $R(i)=2$, then $i+1$ is a trough of the projection vector $V$, and the corresponding trough is $V(i+1)$.
Here we show the results of peaks location search. As shown in Fig. 5, relatively accurate vehicle position points can be obtained after applying the Butterworth low-pass filter to the waterfall diagrams column-wise and subsequently extracting peaks. Through the visual comparison of Fig. 5(a) and (b), it can be seen that if directly extract peaks from denoised signals, it will produce multiple detection phenomenon, which is inconsistent with the actual situation, so it is necessary to carry out Butterworth low-pass filtering.
(a) Wavelet denoised.
(b) Butterworth filtered.
Figure 5: Graphical results of peaks location search algorithm in highway scenario. The blue, orange and green lines represent normalized original signals, wavelet denoised signals and Butterworth filtered signals respectively. The red dots represent the detected peaks.
3.4 Vehicle trajectory extraction
In order to achieve multi-vehicle tracking, we propose a line-by-line matching algorithm based on the vehicle motion model. The vehicle trajectory extraction algorithm flow is shown in Algorithm 2.
@@ 3.4 Vehicle trajectory extraction
2
3Apply Algorithm 1 to $A=(a_{ij})$, returning preprocessed optical fiber matrix $D=(d_{ij})$; vehicle counts $S$ and appearance time of each vehicle $k_{s}$.
4$(k_{s}+\alpha,l_{k_{s}+\alpha})$ is a pair of time and distance position points of each extracted vehicle trajectory;
@@
7
8Select the initial velocity interval $v_{\min}^{(1)}$ to $v_{\max}^{(1)}$;
@@ 9 Calculate the initial distance interval $x_{\min}^{(1)}$ to $x_{\max}^{(1)}$;
10
11foreach $s=1$ to $S$; $\sigma^{(s)}\in\Sigma^{(S)}$ do
@@ 15             Add pair $(k_{s}+1,l_{k_{s}+1})$ to each vehicle trajectory point set $\sigma^{(s)}$;
18             break.
19       end if
20
21 end foreach
22
23foreach $s=1$ to $S$; $\sigma^{(s)}\in\Sigma^{(S)}$ do
@@ 32                   Add pair $(k_{s}+\alpha,l_{k_{s}+\alpha})$ to each vehicle trajectory point set $\sigma^{(s)}$;
Different from previous work [12, 13, 14, 15], complete trajectories of each vehicle in the fixed-size optical fiber diagrams waterfall diagram is regarded as straight lines, that is, the vehicle moves uniformly throughout the whole process. Our algorithm uses this feature locally rather than globally. In each time minimum interval, we view the vehicle as moving in a straight line at uniform speed.
According to Newton’s first law, vehicles can be regarded as moving at a uniform velocity in the absence of lane changes, emergency stops, or overtaking maneuvers. The space-time information reflected by optical fiber within a tiny time unit exhibits a straight-line pattern. This characteristic aligns well with a uniform linear motion model represented as:
$x=vt+x_{0},$ (11)
where $v$ is velocity, $x_{0}$ is the initial position where the vehicle appears, and $x$ is the vehicle’s position at time $t$.
Therefore, after determining the starting position via vehicle position detection, knowing the vehicle’s velocity enables the determination of its position along its trajectory. However, in the case of vehicle congestion, there is a risk of multiple vehicles overlapping. To address this challenge, we expand the speed range of each vehicle from a static line segment to a dynamic line segment. The selection of this dynamic line segment is determined by the slope of the line fitted using historical position and velocity, incorporating a certain degree of confidence. Fig. 6 illustrates the dynamic line segments when vehicle congested, facilitating obtaining key points in vehicle trajectories through line-by-line matching.
Figure 6: Trajectory extraction algorithm. The black line segments represent the distance range and the red dots represent trajectory key points.
$y(x,\omega)=\omega_{0}+\omega_{1}x+\omega_{2}x+\ldots+w_{M}x^{M}=\sum_{j=0}^{M}\omega_{j}x^{j},$ (12)
$E(\omega)=\frac{1}{2}\sum_{n=1}^{N}(y(x_{n},\omega)-t_{n})^{2},$ (13)
Here we illustrate the effectiveness of our proposed tracking method by taking a large truck as an example. Because the truck is too long, two separate trajectories are wrongly detected in the middle sections, as illustrated in Fig. 7 (a). However, our algorithm will detect the key points within the given speed range every second after determining the truck’s entry position. Even if some of the detected key points do not fall on one of the two separate trajectories, it can still reflect the approximate route, which is consistent with the actual situation. Based on key points for a single vehicle, we can deduce the vehicle’s approximate uniform-speed trajectory through polynomial fitting, as shown in Fig. 7 (b).
(a) Trajectory key points.
(b) Trajectory fitting results.
Figure 7: Trajectory extraction results of a large truck. Even if the DAS system detects the truck as two trajectories in the middle sections, our algorithm can still recover the complete trajectory.
4 Experiments
We introduce our self-established highway and tunnel datasets in 4.1. Standard indices regarding experiments in traffic monitoring scenario are illustrated in 4.2. Besides, based on these standard indices, qualitative and quantitative analysis of our comparative experiments are conducted in tunnel and city scenarios in 4.3.
4.1 Datasets
In this section, we conducted data acquisition and analysis in real highway scenario, with a particular focus on a tunnel scenario. The DAS systems are laid 5 to 10cm outside the driveway, with a depth of about 5cm underground. The specific DAS recording parameters are shown in Table 1. As our DAS systems were laid on only one side of the road, the target areas for this paper are explicitly stated. The algorithm proposed aims to extract vehicle trajectories within these target areas. The Jingjintang expressway dataset is used to verify the validity of DAS systems for large-scale traffic monitoring. The Jiurui tunnel dataset and a public city dataset in [12] are used to verify the effectiveness of the proposed algorithm in extracting optical fiber signals. As the description of public city dataset can be found in [12], here we only describe our self-established two datasets.
Table 1: DAS recording parameters.
Parameter Jingjintang expressway dataset Jiurui tunnel dataset
Pulse repetition frequency 4000Hz 4000Hz
Gauge length 300m 5000m
Sample spacing 40cm 80cm
ADA sampling frenquency 4Gbps 4Gbps
@@ 4.1 Datasets
Target area in Jiurui tunnel dataset: The Jiujiang-Ruichang expressway, situated in Jiujiang City, Jiangxi Province, China, represents a real tunnel scenario with a single carriageway featuring two lanes. Our target area focuses on the right lane, disregarding the left lane.
To assess the effectiveness of the optical fiber equipment, cameras were installed on the same right side of the road for data comparison in Jingjintang expressway dataset. Our camera data is used here solely to evaluate the reliability of data collected by optical fiber. The algorithm proposed in this paper exclusively relies on optical fiber data and is independent of camera data.
4.1.1 Jingjintang Expressway Dataset
We selected a 320-meter section of Jingjintang expressway for our study and optical fiber sampling interval is 0.4m. The aerial view of detected area is shown in Fig.8. Simultaneously, a camera mounted on the road captures movement information of the vehicles. Positioned at the starting point of the optical fiber laying side, the camera primarily monitors vehicles in the target area, as depicted in Fig. 9.
Figure 8: Overhead view of the Jingjintang expressway, and the optical fiber with a length of 320m is laid on one side of the road.
@@ 4.1.1 Jingjintang Expressway Dataset
When a vehicle traverses the target area, the DAS systems record corresponding vibration signals. To present the vehicles’ trajectory more intuitively, we utilize waterfall diagrams to plot signal amplitudes, as illustrated in Fig. 10 and Fig. 11.
(a) 3 trucks.
(b) 3 cars.
(c) 1 truck and 4 cars
Figure 10: Simple cases in Jingjintang expressway dataset. Vehicles pass in turn, without interference with each other, and without interference from the opposite lanes.
(a) Traffic flow with cars and trucks.
(b) Multi-vehicle parallel.
(c) Opposite interference.
@@
Figure 11: Complex cases in Jingjintang expressway dataset. (a) Large trucks significantly overlap neighboring small cars. (b) Vehicles pass nearly simultaneously in different lanes. (c) Vehicles from the opposite lane causing interference with vehicles in the target area.
In the waterfall diagrams, the light blue color represents background noise, while prominently colored lines indicate vehicles. The fiber stripe extending from the top left to the bottom right corresponds to vehicles within the target area, and the stripe extending from the top right to the bottom left represents vehicles on the opposite side. Notably, the vehicle’s trajectory is depicted as an approximately straight line within the diagram. This is attributed to the fact that most vehicles on the highway maintain high and consistent velocities when traveling in lanes with unchanged speeds. In this representation, the intercept of the straight line signifies the starting position of the vehicle, and the absolute value of the slope represents its velocity.
Optical fiber waterfall diagrams excel at representing real-time vehicle location data, especially in simple cases where vehicles pass in sequence without interfering with each other, as illustrated in Fig. 10. But during congested period, highway scenario can be more complicated and exist certain inevitable challenges.
@@
(1) Owing to the high sensitivity and rapid response characteristics of optical fiber, large trucks are prone to generating larger signal amplitudes. In instances where a car passes concurrently close, the car’s signals may become entirely submerged beneath the truck’s signals. As shown in Fig. 11(a), it is almost impossible to detect car signals which merge with truck signals.
@@
(2) Considering the relatively large number of vehicles on highways, each vehicle generates substantial signals. When multiple vehicles traverse the highway consecutively, they tend to interfere with each other. As depicted in Fig. 11(b), multiple vehicles overlap with each other and are difficult to distinguish within this cluster.
@@
(3) While the signals from vehicles in the opposite lane are not as clear as those within the target area, signals from trucks in the opposing lane can still interfere with the traffic signal data we aim to capture, as shown in Fig. 11(c).
@@
Additionally, we calculated vehicles to assess the capability of DAS systems. Fig. 12 presents the statistics of vehicle counts for each minute within two hours. Blue lines represent the ground truth counted by the camera, while orange lines represent the fiber stripes counted by the optical fiber. The total number of vehicles within two hours is shown in Table. 2. The optical fiber can capture nearly all the information about trucks with an accuracy rate of 94.88%. Even more accurate to every minute, the fiber stripes can capture almost all large vehicles. However, due to the aforementioned challenges leading to inaccurate optical fiber signals, the counting result for cars is lower than the ground truth only 50%.
(a) Trucks.
(b) Cars.
Figure 12: Vehicle statistics counts in Jingjintang expressway dataset including trucks and cars. The orange and blue lines represent fiber stripes and ground truth respectively.
@@
Table 2: Comparison of real traffic flow and optical fiber stripe counts within 2 hours of Jingjintang expressway dataset.
@@
Vehicle type Fiber stripes Ground truth Accuracy rate
@@
4.1.2 Jiurui Tunnel Dataset
The total length of the Jiurui tunnel is 2340m with DAS systems through the whole tunnel. The aerial view of detected area is shown in Fig.13. The optical fiber sampling interval is 0.8m and duration of data collection is 15min. For subsequent analysis, we selected test sections including an 80-meter section and the entire 2000-meter section. The 80-meter section is used for traffic flow analysis and the 2000-meter section contains special cases such as illegal overtaking and congestion, which are used to verify the effectiveness of the proposed algorithms.
Figure 13: Overhead view of the Jiurui Tunnel, and the optical fiber with a length of 2500m is laid on one side of the road.
Similarly, we draw optical fiber waterfall diagrams to get a more intuitive view of vehicles in Fig. 14. Compared with Jingjintang expressway dataset, the optical fiber stripes in Jiurui tunnel dataset are more strong and obvious. This distinction arises from the tunnel’s configuration, consisting of just two lanes in the same direction, with trucks entering from the outer lane. Consequently, the distinct signals on the waterfall diagram correspond to trucks on the outer lane. Additionally, due to the attenuation effect of the tunnel, the signals from cars in the inner lane are too weak to be adequately recorded. Therefore, the relatively faint signals on the waterfall diagram represent cars in the inner lane.
(a) 2 trucks
(b) 4 trucks
(c) 2 trucks and 4 cars
@@ 4.1.2 Jiurui Tunnel Dataset
The Jiurui tunnel dataset of short distance in Fig. 14 comprises 16 sets of optical fiber data, with each set a collection time of 3-4 minutes and a collection distance of 80 meters. Among these, 7 datasets include corresponding video recordings. Comparing these optical fiber stripes in the optical fiber waterfall diagrams with the actual traffic flow in Fig. 3, due to signal attenuation caused by the tunnel, the fiber stripes of small vehicles on the outer lane were faint, especially electric vehicles, which could not be adequately captured by DAS systems. This explains the observed variance between optical fiber data acquisition and real traffic flow in tunnel scenario.
Table 3: Comparison of real traffic flow and optical fiber stripe counts in Jiurui tunnel dataset of short distance.
Collection time Fiber stripes Ground truth Accuracy rate
@@
Jiurui tunnel long-distance scenario has low-speed limits and restrictions on lane changes and overtaking. Consequently, heavy traffic flows can lead to traffic congestion. Corresponding waterfall diagrams reveal occurrences of multiple lines intersecting (congestion) and even overlapping (illegal overtaking) due to the specific traffic conditions and regulations enforced in the tunnel.
4.2 Metrics
4.2.1 MSE, PSNR and SSIM
To quantify the denoising results, we utilize three indices including mean squared error (MSE), the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM). The formulas are as follows:
${\rm MSE}=\frac{1}{n}\sum\limits_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},$ (14)
${\rm RSNR}=10{\rm log}\frac{v^{2}}{\rm MSE},$ (15)
where $y_{i}$ is the original signals, $\hat{y}_{i}$ is the preprocessed signals, $n$ is total number of sampling points and $v$ is maximum number of image pixels. In general, lower MSE and higher PSNR values imply better reconstruction performance. SSIM is a full-reference evaluation index considering the brightness, contrast and structure of images [37]. The formulas are as follows:
${\rm SSIM}=[l(y,\hat{y})]^{\alpha}[c(y,\hat{y})]^{\beta}[s(y,\hat{y})]^{\gamma},$ (16)
where $l(y,\hat{y}),c(y,\hat{y}),s(y,\hat{y})$ stand for luminance, contrast and structure respectively. SSIM value ranges from -1 to 1. The closer it is to 1, the better the denoising effect is.
4.2.2 Traffic indices